Zach and Koji
The old small MMT was removed and wrapped by Al foils.
The steering mirror IM2-IM4 were displaced and aligned.
The Faraday isolator block is moved and aligned.
The MC is realigned and resonatng TEM-00.
Now the MC has slightly miscentered beam on the mirrors owing to change of the stack leveling.
OSEMs are also in a strange state. We should check this later.